What concentration of AlCl3 will produce a pH of 2.50 in 500ml of H2O
1
Expert's answer
2020-06-04T10:37:08-0400

Solution:

AlCl3=Al3++3Cl- - Aluminum chloride completely dissociates in an aqueous solution.

Al3++H2O=AlOH2++H+ - only the first stage of hydrolysis proceeds significantly.

Kh=[AlOH2+][H+]/[Al3+] - constant of the first stage of Al3+ ion hydrolysis.

Multiplying the numerator and denominator of the hydrolysis constant expression by [OH-], we get:

Kh=[AlOH2+][H+][OH-]/([Al3+][OH-])=Kw/KbAlOH2+

Kw=[H+][OH-]=1*10-14 - Ionic product of water

KbAlOH2+=[Al3+][OH-]/[AlOH2+]=1*10-9 - the constant of basicity of the AlOH2+ ion

For a dilute solution it will be fair:

[AlOH2+]=[H+]

and [Al3+]=CAlCl3

Therefore

Kw/KbAlOH2+=[H+]2/CAlCl3

CAlCl3=[H+]2KbAlOH2+/Kw

pH=-lg[H+]

[H+]=10-pH

So, CAlCl3=10-2pHKbAlOH2+/Kw

CAlCl3=10-2*2.5*1*10-9/1*10-14=1(mol/L)

Answer:

Concentration of AlCl3 that will produce a pH of 2.50 in 500ml of H2O is 1mol/L.

And 0.5 mol of AlCl3 should be dissolved in 500ml of an aqueous solution.